# Bayesian Optimization in Language space: An Eval-Efficient AI Self-Improvement Framework

### v1

@@ Abstract
Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refine its own outcomes. Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement-learning fine-tuned LLMs. Here, their ‘performance’ is typically measured by query efficiency—the number of LLM-generated solution samples required to reach a specified performance threshold. However, in many societal applications, the primary limitation is not generating new solutions but evaluating them. For instance, evaluating an ad’s effectiveness requires significant human feedback, which is far more costly and time-consuming than generating a candidate ad. To optimize for the evaluation efficiency objective, a natural approach is to extend Bayesian Optimization (BO), a framework proven optimal for evaluation efficiency, to the language domain. However, the difficulty of directly estimating suitable acquisition functions in LLMs’ minds makes this extension challenging. This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model) statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency. Based on this result, we propose TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), a simple and eval-efficient language-space Bayesian optimization framework for AI self-improvement. We also empirically validate T-BoN BO by applying it to automated ad alignment tasks for persona distribution, demonstrating its superior performance compared to popular state-of-the-art baselines. In sum, T-BoN BO is a simple but principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
@@ 1.1 Self-improving AI
Data-centric decision making has long relied on a human-driven cycle of historical data analysis, solution proposal, and experimentation. This human-driven iterative approach has supported a wide range of business decisions, ranging from website design (hauser2009website) to product recommendations (nandy2021b). We briefly describe this paradigm using the example of ad-optimization; see the left panel of Figure 1. The process starts with a manager generating a set of ad creatives based on a creative brief. This task is often done by ad agencies or marketers, who rely on their understanding of the product, theories of consumer behavior, market conditions, and feedback from prior ad campaigns. This ad generation process typically takes weeks, and serves as the main bottleneck in the iterative ad-optimization process. Once we have a set of candidate ads, the manager moves to evaluation, i.e., feeds these ads into advertising platforms or consumer surveys to identify the best performing one, and this step usually takes a few days. Finally, to complete this feedback loop, the manager analyzes the performance of the ads from this campaign and uses the insights learned to generate a new slate of candidate ads; and the iterative process continues.
Although effective in many settings, this human-driven workflow suffers from drawbacks. First, the speed with which firms can iterate over this process depends critically on how quickly humans can learn from the performance of prior solutions and propose new candidate solutions. Historically, generating new candidate solutions (or ads in the case of our example) has been time-consuming, and this step formed a key bottleneck in speeding up the process. Further, there exists no formal framework for how human agents can learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition. Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024); but how to leverage insights from prior rounds of evaluation systematically and improve the set of candidate solutions remains an unsolved problem.
Ad generation (Step 1)(Weeks)Analysis (Step 3)(days)Ad evaluation (Step 2)(days)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(days)AIBottleneck
@@
Recent advances in generative AI have begun to transform this human-driven paradigm by providing potential solutions to the issues discussed above. First, candidate generation is fast and scalable with generative AI models. For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power) (see right panel of Figure 1). [^1] Indeed, a substantial share of creatives is now generated by AI, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing). Second, Large language Models (LLMs) and related generative systems can now conduct sophisticated data analysis, summarizing performance patterns and identifying potential drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners), providing a pathway for analyzing historical data and improving candidate solutions. As such, with the new capabilities enabled by AI, the traditional bottlenecks of analysis and generation of candidate solutions are significantly reduced. Given these developments, there is immense interest in the development and deployment of self-improving AI (silver2025welcome), which automates the complete analysis–generation–evaluation loop at scale, allowing systems to update themselves continuously in response to emerging data (MantiaChatterjeeLee2025; chen2025xbench).
@@ 1.2 Research Agenda and Challenges
In this paper, we study a prompt-optimization-based self-improving AI framework for societal and business problems, motivated by recent successes of prompt-optimization-based self-improving AI algorithms that outperform state-of-the-art reinforcement learning fine-tuning methods (agrawal2025gepa). While the use of modern generative AI systems has reduced the overhead associated with content generation and analysis (Steps 1 and 3), there are still three challenges that we need to overcome to design an efficient and effective self-improving AI that can automate decision-making in societal and business problems.
@@
    Challenge 1 – Evaluation-efficiency
    First, in many real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (query efficiency) but evaluating them efficiently (evaluation efficiency). For instance, in the case of digital advertising, generating ad creatives using LLM queries is super fast, but evaluating an ad’s effectiveness requires significant human feedback or large-scale experimentation (johnson2023inferno). As such, any AI system that we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than trying to optimize for a few inexpensive generations as possible. While there is a nascent but growing literature in computer science on self-improving AI systems, this literature mainly focuses on generation efficiency as the main bottleneck (agrawal2025gepa), and has largely overlooked evaluation efficiency.
@@
    Challenge 2 – Lack of gradient functions
    Second, notice that we are essentially solving an optimization problem in language space, i.e., we want to take an input prompt, evaluate its performance, and improve it to reach a higher performance. [^2] Self-improving AI can thus be viewed as principled prompt optimization or context engineering since it involves improving textual inputs/prompts based on historic performance data (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro). In standard numerical settings, the natural way to solve such problems is to use gradient functions. However, in language spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Thus, we first need to define a meaningful notion of a “gradient” in the domain of text.
  • $\bullet$
    Challenge 3 – Balancing exploration and exploitation
    Third, while it is tempting to move fully along the textual gradient that worked well previously, i.e., exploit the knowledge gained from prior prompts, this can be sub-optimal due to uncertainty about that knowledge. Thus, we also need to bake some exploration in the choice of gradient direction. Again, while there exist bandit-based approaches that balance exploration and exploitation in the case where we have a fixed candidate set and the exploration space is numerical (jamieson2014best; russo2016simple; russo2018tutorial; mason2020finding; simchi2023multi; fiez_etal_2024), these do not automatically translate to our setting since our goal is to explore in the language space and our exploration focuses on changing the set of candidate solutions rather than deciding how much to explore each option within a fixed candidate set. Thus, a key challenge is to account for uncertainty to guide the search over the language space. Canonical uncertainty measures, such as the Upper Confidence Bound (UCB) or the posterior distribution, are not readily available from off-the-shelf LLMs, making their direct application to language space non-trivial. Further, we note that self-consistency variance or embedding distances are poor proxies, and training a Gaussian Process over high-dimensional text embeddings is brittle, difficult to implement, expensive, and data-hungry.
@@
In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that effectively balances exploration and exploitation in the language space.
1.3 Our Framework and Algorithm
We start with the observation that prompt optimization for self-improving AI follows a familiar pattern – we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next while minimizing evaluations. Throughout, we have to be cognizant of three concerns: (1) we need to balance exploration and exploitation when choosing the next candidate solution, (2) we have no closed-form model linking prompts to scores, and (3) the main scarce resource is the number of evaluations rather than candidate generation. Our key insight is that these features of our problem are similar to the regime in which classic Bayesian Optimization (BO) operates successfully (frazier2018tutorial; srinivas2012information; papenmeier2025understanding). BO is a principled framework for tackling black-box optimization problems where evaluation efficiency is an important consideration. (See Web Appendix E for a detailed theoretical discussion of BO.) Thus, we seek to extend BO from numerical inputs to the space of natural language prompts to design an evaluation-efficient self-improving AI.
At its core, BO assumes a surrogate model (e.g., Gaussian Process or Reproducing Hilbert Kernel Space) to approximate the objective function and provide uncertainty estimates of that approximation. From this surrogate, BO constructs a function often called the acquisition function and selects as the next candidate the point with the highest acquisition value. One of the most widely used BO methods is Bayesian Optimization with Upper Confidence Bound (UCB) acquisition function (srinivas2012information). In practice, since the global maximizer of the UCB acquisition function is often not available in a closed form, we typically use gradient-based BO with UCB acquisition function (wilson2018maximizing). Gradient-based BO methods have shown surprising effectiveness for high-dimensional problems when the gradient-based BO is adapted to avoid vanishing gradients, for example, through random multi-start (where each initialization yields a local maximum of UCB) and careful gradient sampling strategies (papenmeier2025understanding). However, the standard gradient-based UCB BO is not directly applicable to the language space, as the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Therefore, we first need to define a notion of gradients in the language space and devise an appropriate acquisition function in the text space.
We now describe how we modify and adapt the BO framework to address the challenges discussed earlier. We propose a simple, evaluation-efficient iterative prompt-optimization framework that we call TextGrad-Best-of-N Bayesian Optimization (T-BoN BO) that addresses all three challenges discussed earlier. First, the BO framework naturally addresses the first challenge of the evaluation efficiency. It is known that BO with UCB acquisition function (srinivas2012information) is a theoretically evaluation-efficient algorithm, i.e., optimally balances exploration and exploitation, minimizing regret (whitehouse2023sublinear; wang2023regret). Specifically, for both expected cumulative regret and simple regret, the regret upper bound of BO with UCB acquisition function tightly matches the regret lower bound of scarlett2017lower.
@@ 1.3 Our Framework and Algorithm
To solve the second challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapt gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space. This makes gradient-style search feasible without continuous relaxations or mapping the LLM prompts to the embedding space. textgrad show that textual gradients based on their approach enable consistent improvements across diverse scientific domains such as code-completion benchmarks, improving accuracy on graduate-level QA datasets, optimizing drug-like molecules, and refining radiotherapy planning for clinical objectives. These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in language space.
@@
To address the third challenge, T-BoN BO chooses the next prompt to evaluate via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling). Moving in the textual gradient direction suggested by the critic exploits the expected improvement implied by past results but ignores uncertainty in those expectations. T-BoN bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of the optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate. We show that Best-of-$N$ over locally sampled textual edits statistically emulates gradient ascent on a UCB acquisition function, so that Best-of-$N$ effectively acts as an implicit optimistic acquisition maximizer in language space.
@@
We now briefly describe our T-BoN BO algorithm. At each optimization iteration, in Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly. In Step 2 (Best-of-$N$ textual-gradient phase), we perform a sequence of textual gradient updates: for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt. After iterating Step 2, we obtain a refined prompt and its associated solution to evaluate. In Step 3 (evaluation phase), we evaluate this solution to obtain a scalar score, which is then used by T-BoN BO to update its history and meta-reflection for subsequent optimization steps.

### v2

@@ Abstract
Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refines its own outcomes. Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement learning fine-tuned LLMs. Here, their ‘performance’ is typically measured by query efficiency—the number of LLM-generated solution samples required to reach a specified performance threshold. However, in many societal applications, the primary limitation is not generating new solutions but evaluating them. For instance, evaluating an ad’s effectiveness requires significant human feedback, which is far more costly and time-consuming than generating a candidate ad. To optimize for the evaluation efficiency objective, a natural approach is to extend Bayesian Optimization (BO), a framework proven optimal for evaluation efficiency, to the language domain. However, the difficulty of directly estimating suitable acquisition functions in LLMs’ minds makes this extension challenging. This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model), statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency. Based on this result, we propose TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), a simple and eval-efficient language-space Bayesian optimization framework for AI self-improvement. We also empirically validate T-BoN BO by applying it to automated ad alignment tasks for persona distribution, demonstrating its superior performance compared to popular state-of-the-art baselines. In sum, T-BoN BO is a simple but principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
@@ 1.1 Self-improving AI
Data-centric decision making has long relied on a human-driven cycle of historical data analysis, solution proposal, and experimentation. This human-driven iterative approach has supported a wide range of business decisions, ranging from website design (hauser2009website) to product recommendations (nandy2021b). We briefly describe this paradigm using the example of ad-optimization; see the left panel of Figure 1. The process starts with generating a set of ad creatives based on a creative brief. This task is often done by ad agencies or marketers, who rely on their understanding of the product, theories of consumer behavior, market conditions, and feedback from prior ad campaigns. This ad generation process typically takes weeks, and serves as the main bottleneck in the iterative ad-optimization process. Once we have a set of candidate ads, the manager moves to evaluation, i.e., feeds these ads into advertising platforms or consumer surveys to identify the best performing one, and this step usually takes a few days. Finally, to complete this feedback loop, the manager analyzes the performance of the ads from this campaign and uses the insights learned to generate a new slate of candidate ads; and the iterative process continues.
Although effective in many settings, this human-driven workflow suffers from drawbacks. First, the speed with which firms can iterate over this process depends critically on how quickly humans can generate new candidate solutions. Historically, generating new candidate solutions (or ads in the case of our example) has been time-consuming, and this step formed a key bottleneck in speeding up the process. Further, there exists no formal framework for how human agents can analyze and learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition. Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024).
Ad generation (Step 1)(Weeks)Analysis (Step 3)(Days)Ad evaluation (Step 2)(Days-weeks)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(Days-weeks)AIBottleneck
@@
Recent advances in generative AI have begun to transform this human-driven paradigm by providing potential solutions to the issues discussed above. First, candidate generation is fast and scalable with generative AI models. For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power). [^1] Indeed, a substantial share of creatives is now generated by AI, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing). Second, Large language Models (LLMs) and related generative systems can now conduct sophisticated data analysis, summarizing performance patterns and identifying potential drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners), providing a pathway for analyzing historical data and improving candidate solutions. As such, with the new capabilities enabled by AI, the time required for previous data analysis and generation of new candidate solutions are significantly reduced.
Given these developments, there is immense interest in the development and deployment of AI-based automation of the complete generation–evaluation-analysis loop at scale, allowing systems to update themselves continuously in response to emerging data (See right panel of Figure 1). In this loop, generative models produce new solution candidates, data-science agents conduct evaluation of the solution candidates, and multi-modal reasoning LLM-based analytical agents interpret the evaluation results and synthesize insights to guide the next generation step. This type of automated cycle running system is often referred to as the self-improving AI or self-evolving AI (MantiaChatterjeeLee2025; chen2025xbench; silver2025welcome).
This transformation has an important implication for practice: it dramatically lowers the marginal cost of iteration, making continuous experimentation feasible not only for large firms or organizations with dedicated analytics teams but also for small and medium-sized firms that historically lacked the creative and analytical capacity to run many cycles in parallel. When generation and analysis become cheap and largely automated, even resource-constrained organizations can maintain always-on optimization loops that adapt campaigns to evolving market conditions and consumer responses.
As generation and analysis become increasingly automated and inexpensive in this self-improving loop, the bottleneck in data-centric decision making shifts from proposing and interpreting candidates to evaluating them in the real world. Even if analysis can be performed quickly and new candidate solutions can be generated almost instantaneously, each iteration still has to be validated by interacting with the real environment and collecting outcome data. The effective pace of the cycle is therefore governed by evaluation-related constraints such as available traffic, experimentation budgets, and the amount of data required to obtain statistically reliable comparisons among candidates, rather than by generation or analysis speed.
For instance, in digital advertising, even when generating ad creatives via LLM queries is super fast, but evaluating an ad’s effectiveness still requires significant human feedback or large-scale experimentation (johnson2023inferno). A similar pattern arises in other scientific and engineering domains: protein structures or sequences can be proposed rapidly by models such as AlphaFold (jumper2021highly), but their functional properties must still be confirmed through slow and noisy wet-lab assays; and chip layouts can be generated in silico by systems such as AlphaChip (goldie2024chip), yet rigorous evaluation depends on physical fabrication and benchmarking. As generation and analysis become increasingly automated and inexpensive, the evaluation step thus emerges as the primary bottleneck that limits the overall cadence of self-improvement. As such, any societal and business AI system we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than optimizing for a few inexpensive generations.
1.2 Prompt optimization-based self-improving AI
A natural question is, then, how the computer science literature has instantiated self-improving AI in concrete algorithmic frameworks. To date, a central line of success in self-improving AI has arisen in the domain of iterative prompt optimization, where AI agents iteratively refine their own prompts in response to performance feedback. agrawal2025gepa showed that iterative prompt optimization-based self-improving AI can be much more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek). Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments. Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage.
In all those benchmarks, the primary efficiency metric was generation efficiency. i.e., the number of LLM query calls required to generate candidate answers to reach a target performance level. Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, not evaluation efficiency. Indeed, leading approaches are often specified as procedural heuristic search routines which finds near-optimal solution with small computations, e.g., population-based or random-search-style algorithms that treat prompts as individuals in an evolutionary process or as configurations in a heuristic pipeline. Specifically, they often maintain a pool or batch of prompts, repeatedly generate new variants via LLM-driven mutation, crossover, reflection, or trace-based resampling, score these candidates on held-out data or rollouts, retain the top-performing or Pareto-optimal prompts, and iterate this generate–score–select–mutate loop under a budget of LLM forward passes (10.5555/3692070.3692611; guo2024connecting; agrawal2025gepa; better_together). This mirrors the design philosophy of classical evolutionary and heuristic optimization, where the quality of the best solution is assessed as a function of the computation budget rather than the evaluation budget (nguyen2016understanding; ding2023quality; hu2024automated; liu2024evolution). Consequently, evaluation efficiency objective has been largely overlooked in the computer science literature.
However, as we discussed earlier, in self-improving AI for real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (generation efficiency) but evaluating them efficiently (evaluation efficiency). This shift in the dominant constraint calls for rethinking prompt-optimization-based self-improving AI frameworks so that they explicitly treat evaluation budget as the scarce resource to be allocated and optimized over.
@@ 1.3 Research agenda and Challenges
Motivated by the effectiveness of iterative prompt-optimization-based self-improving AI and the overlooked importance of evaluation efficiency in existing frameworks, we now turn to our central question: how can we design a prompt-optimization-based self-improving AI framework that is evaluation-efficient? In addition to evaluation efficiency, we would like the framework to be easy to implement, theoretically grounded, and empirically perform well. However, realizing these desiderata in a unified framework is nontrivial.
The key insight in this paper is that prompt-optimization-based self-improving AI follows a familiar pattern we saw in the optimization framework called Bayesian Optimization (BO) (frazier2018tutorial; srinivas2012information; papenmeier2025understanding). This insight is grounded in three observations: First, LLMs learn in-context from previous evaluations (brown2020language; coda2023meta; lu2024emergent; cahyawijaya2024llms), acting as Bayesian learners (xie2021explanation; wang2023large; wakayama2025context). Second, we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next. Third, we have to be cognizant of three concerns: (1) we have no closed-form model linking prompts to scores, (2) we need to balance exploration and exploitation when choosing the next candidate solution, and (3) the goal is to optimize for the number of evaluations rather than candidate generation.
BO is a principled framework for tackling such black-box optimization problems with such features. In BO, we build a function called the acquisition function and choose the point with the top of the acquisition function as the next point to evaluate. Among many variants of BO, Upper Confidence Bound (UCB)-BO (srinivas2012information), a version of BO with UCB acquisition function, is proven to be evaluation-efficient, in the sense that no algorithm can achieve better regret in terms of both simple regret and expected cumulative regret (whitehouse2023sublinear; wang2023regret). Figure 2 illustrates the procedure of UCB-BO.
Figure 2: An example of a non-parallel UCB-BO procedure for four optimization steps. At each step, we choose the maximum (or multiple local maxima in the parallel UCB-BO) of the blue UCB function as the next point to evaluate.
As you can see from Figure 2, the UCB acquisition function is often non-concave. To resolve this issue, in practice, we often use parallel gradient-based UCB-BO (wilson2018maximizing), where we 1) start from multiple initial points and take gradients to find a set of local optima, and 2) choose those local optima as the points to evaluate next. Parallel gradient-based BO has shown surprising empirical effectiveness for high-dimensional problems (papenmeier2025understanding).
Building on the key insight, we propose to achieve evaluation efficiency by casting iterative prompt optimization-based self-improving AI as parallel gradient-based UCB Bayesian optimization in language space. At each iteration, we treat the prompts from the previous round as initialization points and take gradient steps to ascend the UCB acquisition function toward its local optima. We then select which prompts to evaluate from among these local optima. After each evaluation round, the LLM and its associated critic models implicitly update a surrogate belief over prompt quality via in-context learning and corresponding UCB, analogous to the posterior update in classical Bayesian optimization.
However, there are two challenges we need to overcome.
@@
    Challenge 1– Lack of gradient functions
    In prompt spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Thus, we first need to define a meaningful notion of a “gradient” in the domain of text. Although one might consider going around this challenge by embedding the prompt first, taking the gradient there, and reverting the prompt, this strategy runs into the well-known soft-to-discrete projection problem, where such reverted prompts break any clean correspondence between infinitesimal moves in embedding space and interpretable edits of the underlying natural-language prompt (cui2025automatic).
@@
    Challenge 2 – Taking gradients for UCB
    Second, even if we can define meaningful gradients in prompt space, we do not have direct access to the underlying UCB acquisition function whose gradient we would like to follow. The surrogate posterior over prompt quality that induces UCB lives implicitly in the LLM, rather than as an explicit Gaussian-process-style model that exposes closed-form posterior means, variances, and their derivatives. Consequently, standard gradient-based BO techniques that differentiate the UCB constructed from a known surrogate cannot be applied naively: we must design a mechanism that elicits UCB-like, optimism-in-the-face-of-uncertainty behavior using only black-box text interactions and critic feedback, without ever computing the UCB function or its gradient in closed form.
@@
In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that optimally balances exploration and exploitation in the language space.
1.4 Proposed Method
Figure 3: Gradient steps of parallel gradient-based BO (Left image) and those of T-BoN BO (Right image). T-BoN BO extends parallel gradient-based BO to the language space.
We propose a simple, evaluation-efficient iterative prompt-optimization framework, called TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), that adapts parallel gradient-based UCB-BO for prompt optimization based self-improving AI (Figure 3), addressing both of the challenges discussed earlier.
@@ 1.4 Proposed Method
To solve the first challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapts gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space. This makes gradient-style search feasible without continuous relaxations or mapping the LLM prompts to the embedding space. For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal. textgrad show that textual gradients based on their approach enable consistent improvements across diverse scientific domains such as code-completion benchmarks, improving accuracy on graduate-level QA datasets, optimizing drug-like molecules, and refining radiotherapy planning for clinical objectives. These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in language space.
@@
To address the second challenge, T-BoN BO chooses the prompt to evaluate next via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling), resulting in the gradient direction we call Best-of-N gradient. Moving in the textual gradient direction suggested by the critic exploits the expected improvement implied by past results but ignores uncertainty in those expectations. Therefore, it does not climb the UCB acquisition function. Best-of-N gradient bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of UCB’s optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate. That is, the Best-of-$N$ gradient is designed to simulate the implicit UCB acquisition gradient in language space.
@@
We now briefly describe our T-BoN BO algorithm. At each optimization iteration, in Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly. In Step 2 (Best-of-$N$ textual-gradient phase), we perform a sequence of Best-of-N gradient updates from the current prompt. That is, for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt. After iterating Step 2, we obtain a refined prompt and its associated solution to evaluate. In Step 3 (evaluation phase), we evaluate this solution to get a scalar score, which T-BoN BO then uses to update its history and meta-reflection for subsequent optimization steps.